# Data-driven algorithm design using neural networks with applications to branch-and-cut

### v1

@@ 3.1 Preliminaries
The most successful algorithms and solvers for integer programming problems are based on a methodology called branch-and-cut. In a branch-and-cut algorithm, one maintains two things in every iteration of the algorithm:
  1. 1.
    a current guess for the optimal solution,
  2. 2.
    a collection of polyhedra that are subsets of the original polyhedral relaxation of the ILP.
In every iteration, one of these polyhedra are selected and the continuous linear programming (LP) solution for that selected polyhedron is computed. If the solution has objective value worse than the current guess, this polyhedron is discarded from the list and the algorithm moves to the next iteration. Otherwise, if the solution is integral, the guess is updated with this integral solution and this polyhedron is removed from further consideration. If the LP solution is not integral, one decides to either add some cutting planes or branch. In the former case, additional linear constraints are added to this polyhedron under consideration without eliminating any feasible solutions. In the latter case, one selects a fractional variable $\mathbf{x}_{i}$ in the LP solution and partitions the current polyhedron into two polyhedra by adding constraints $\mathbf{x}_{i}\leq\lfloor f_{i}\rfloor$ and $\mathbf{x}_{i}\geq\lfloor f_{i}\rfloor+1$, where $f_{i}$ is the value of this fractional variable. The current polyhedron is then replaced in the list by these two new polyhedra. This entire process can be tracked by a branch-and-cut tree whose nodes are precisely the different polyhedra processed by the algorithm. The algorithm terminates when there are no more polyhedra left in the active list and the current guess is reported as the optimal solution. As is often done in practice, an a priori bound $B\in\mathbb{Z}_{+}$ is set on the size of a tree; if this bound is exceeded by the algorithm at any stage, the algorithm exist early and the current guess for the solution is returned. The branch-and-cut tree size is a very good indication of how long the algorithm takes to solve the problem since the main time is spent on solving the individual LPs in the iterations of the algorithm. We will thus use the tree size as the “score" function to decide how well branch-and-cut did on any instance.
There are many different strategies to generate cutting planes in branch-and-cut (Conforti et al., 2014; Nemhauser and Wolsey, 1988; Schrijver, 1986). We will focus on the so-called Chvátal-Gomory (CG) cutting planes and Gomory Mixed-Integer (GMI) cuts (Conforti et al., 2014). There are usually several choices of such cutting planes to add (and some families are even infinite in size (Conforti et al., 2014)). We wish to apply the results of Section 2 to decide which cutting plane to select so that the branch-and-cut tree size is small.
@@ 3.2 Learnability of parameterized CG cut(s)
Following the discussion in (Balcan et al., 2021b), we define $f_{\operatorname{CG}}(I,\mathbf{u})$ as the size of the branch-and-bound tree for a given ILP instance $I\in\mathcal{I}$ with a CG cut parameterized by a multiplier $\mathbf{u}\in[0,1]^{m}$ added at the root. We interpret $f_{\operatorname{CG}}$ as a score function elaborated in Section 2.2. The piecewise structure of $f_{\operatorname{CG}}$ in its parameters is characterized by:
Lemma 3.2 (Lemma 3.2 in (Balcan et al., 2021b)).
@@
Theorem 3.3.
Under the same conditions as Theorem 2.5 and 2.6, with the score function $f_{\operatorname{CG}}$, we have
@@
Lemma 3.4 (Lemma 3.4 in (Balcan et al., 2021b)).
@@
Theorem 3.5.
Under the same conditions as Theorem 2.5 and 2.6, with the score function $f_{\operatorname{CG}}^{k}$, we have
@@ 3.3 Learnability of cutting plane(s) from a finite set
The selection of an optimal cut from an infinite pool of candidate cuts, as discussed in Theorem 3.3 and Theorem 3.5, is often difficult and inefficient in practice. Consequently, a popular way is to select cuts based on information from the simplex tableau (such as GMI cuts), as well as some combinatorial cuts, which inherently limit the number of candidate cuts considered to be finite.
Suppose we have $r$ candidate cuts $\mathcal{C}=\{c_{1},\dots,c_{r}\}$, and we define $f_{\operatorname{ROW}}(I,c)$ as the branch-and-bound tree size after adding a cut $c\in\mathcal{C}$ at the root for a given ILP instance $I\in\mathcal{I}$. Then Corollary 2.7 implies the following theorem:
Theorem 3.6.
Under the same conditions as Theorem 2.5 and 2.6, with the score function $f_{\operatorname{ROW}}$, we have
$\operatorname{Pdim}\left(\mathcal{F}_{N^{\operatorname{sgn}},\operatorname{Sigmoid}}^{f_{\operatorname{ROW}}}\right)$ $=\mathcal{O}\left(W\log(Ur)\right),$
$\operatorname{Pdim}\left(\mathcal{F}_{N^{\operatorname{ReLU}},\operatorname{CReLU}}^{f_{\operatorname{ROW}}}\right)$ $=\mathcal{O}\left(LW\log(U+r)\right).$
4 Experiments
@@
This problem, concerning the selection from an infinite pool of cuts, presents a significant challenge. The target function $f_{\operatorname{CG}}$, in relation to its parameters, is an intricately complex, high-dimensional, and piecewise constant function with numerous pieces (recall Lemma 3.2), making direct application of classical gradient-based methods seemingly impractical. As such, we use a reinforcement learning (RL) approach, with our goal switched to identify a relatively “good” neural network for this distribution.
@@ 4.1 Experimental setup
We consider the multiple knapsack problems (Kellerer et al., 2004) with 16 items and 2 knapsacks, using the Chvátal distribution as utilized in (Balcan et al., 2021c), originally proposed by Chvátal 1980. Our synthetic dataset has a training set of 5,000 instances and a test set of 1,000 instances from the same distribution.
@@
Each instance $I$ sampled from the distribution $\mathcal{D}$ is treated as a state in RL, with the outputs of the neural network considered as actions in the RL framework. The neural network thus functions as the actor in an actor-critic scheme (Silver et al., 2014), where the reward is defined as the percentage reduction in the tree size after adding a cut, i.e., $\frac{f_{\operatorname{CG}}(I,\mathbf{0})-f_{\operatorname{CG}}(I,\mathbf{u})}{f_{\operatorname{CG}}(I,\mathbf{0})}$. The Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm (Fujimoto et al., 2018) is used here for the training of the neural network.
All experiments are conducted on a machine with an Intel Core i7-12700F 12-Core CPU and an RTX 3070 GPU. Gurobi 10.0.3 (Gurobi Optimization, LLC, 2023) serves as the solver for the ILPs, with the neural networks being implemented using PyTorch 2.0.1.
@@ 4.2 Empirical results
The experimental results indicate that even a suboptimal neural network parameterization can outperform the cut selection methodologies used in state-of-the-art solvers like SCIP (Gamrath et al., 2020), which relies on the combination of some auxiliary score functions
$\mathbf{u}^{*}\in\mathop{\operatorname{arg\,max}}\limits_{\mathbf{u}\in\mathcal{U}}\sum_{i=1}^{\ell}\mu_{i}\operatorname{Score}_{i}(\mathbf{u},I),$ (3)
where $\operatorname{Score}_{i}$ represents a range of heuristic scoring functions that evaluate various aspects of a cut, such as its “Efficacy(Balas et al., 1996) and “Parallelism(Achterberg, 2007), for a specific instance $I\in\mathcal{I}$. The coefficients $\mu_{i}$ serve as tunable weights for these scoring models. The set $\mathcal{U}$, including all the potential cut parameters to be considered, is finite when limited to CG cuts derived from the simplex tableau.
Figure 1: Comparison of branch-and-bound tree sizes using different CG cut selection strategies.
Figure 1 presents the average tree size comparison across 1,000 novel instances, with four distinct strategies: 1) The blue solid line represents the tree size when employing CG cuts selected from the simplex tableau based on the highest combined score of cut Efficacy and Parallelism, modulated by a parameter $\mu$ (incremented in steps of 0.01); 2) The red dotted line indicates larger tree sizes resulting from a randomized selection of CG cuts from the tableau; 3) The green dash-dotted line demonstrates a notable reduction in tree size when incorporating cuts generated through our RL approach; 4) The purple dash-dotted line follows the same approach as 3), but uses linear threshold neural networks to generate CG cut parameters, backpropagated using the idea of the Straight-Through Estimator (Bengio et al., 2013).
@@
Table 1: Comparison of the computational speeds between neural network inference and LP solving. This table presents the total time in seconds for 100 runs on a test set of 1,000 instances.
@@
Computing cuts via trained neural network on GPU 0.030
Computing cuts via trained neural network on CPU 0.139
Solving required LPs using Gurobi 4.219
@@ 5 Discussions and open questions
In our study, we concentrated on adding CG cuts solely at the root of the branch-and-cut tree. However, devising a strategy that generates high quality cutting planes while being efficient across the entire branch-and-cut tree poses a significant and intriguing challenge for future research. Further, our theoretical findings are applicable to any encoder that maps instances into Euclidean spaces. Hence, utilizing a fixed, pretrained encoder capable of converting ILPs with different number of constraints into a same Euclidean space can in principle enable the training of a unified neural network to generate cutting planes across the branch-and-cut tree. Moreover, an effective encoder could improve the neural network’s performance beyond that achieved with the basic stacking encoder used in our paper.
The neural network training problem (2) also requires further study. We used RL to update neural network parameters and minimize average tree size for ILP instances from a given distribution. However, this method does not guarantee convergence to optimal neural parameters and relies heavily on the exploratory nature of the RL algorithm. For ILP distributions where random exploration of Chvátal multipliers is unlikely to yield smaller tree sizes, the RL algorithm may struggle to identify an effective parameter setting. Developing a more efficient and robust training methodology would greatly improve the practical value of our work.
Acknowledgements
@@
Appendix A Proofs of Main Results

### v2

@@ 3.1 Preliminaries
The most successful algorithms and solvers for integer programming problems are based on a methodology called branch-and-cut. In a branch-and-cut algorithm, one maintains two things in every iteration of the algorithm: 1) a current guess for the optimal solution, 2) a collection of polyhedra that are subsets of the original polyhedral relaxation of the ILP. In every iteration, one of these polyhedra are selected and the continuous linear programming (LP) solution for that selected polyhedron is computed. If the solution has objective value worse than the current guess, this polyhedron is discarded from the list and the algorithm moves to the next iteration. Otherwise, if the solution is integral, the guess is updated with this integral solution and this polyhedron is removed from further consideration. If the LP solution is not integral, one decides to either add some cutting planes or branch. In the former case, additional linear constraints are added to this polyhedron under consideration without eliminating any feasible solutions. In the latter case, one selects a fractional variable $\mathbf{x}_{i}$ in the LP solution and partitions the current polyhedron into two polyhedra by adding constraints $\mathbf{x}_{i}\leq\lfloor f_{i}\rfloor$ and $\mathbf{x}_{i}\geq\lfloor f_{i}\rfloor+1$, where $f_{i}$ is the value of this fractional variable. The current polyhedron is then replaced in the list by these two new polyhedra. This entire process can be tracked by a branch-and-cut tree whose nodes are precisely the different polyhedra processed by the algorithm. The algorithm terminates when there are no more polyhedra left in the active list and the current guess is reported as the optimal solution. As is often done in practice, an a priori bound $B$ is set on the size of a tree; if this bound is exceeded by the algorithm at any stage, the algorithm exist early and the current guess for the solution is returned.
The branch-and-cut tree size is a very good indication of how long the algorithm takes to solve the problem since the main time is spent on solving the individual LPs in the iterations of the algorithm. We will thus use the tree size as the “score" function to decide how well branch-and-cut did on any instance.
There are many different strategies to generate cutting planes in branch-and-cut Conforti et al. (2014); Nemhauser and Wolsey (1988); Schrijver (1986). We will focus on the so-called Chvátal-Gomory (CG) cutting planes and Gomory Mixed-Integer (GMI) cuts Conforti et al. (2014). There are usually several choices of such cutting planes to add (and some families are even infinite in size Conforti et al. (2014)). We wish to apply the results of Section 2 to decide which cutting plane to select so that the branch-and-cut tree size is small.
@@ 3.2 Learnability of parameterized CG cut(s)
Following the discussion in Balcan et al. (2021d), we define $f_{\operatorname{CG}}(I,\mathbf{u})$ as the size of the branch-and-bound tree for a given ILP instance $I\in\mathcal{I}$ with a CG cut parameterized by a multiplier $\mathbf{u}\in[0,1]^{m}$ added at the root. We interpret $f_{\operatorname{CG}}$ as a score function elaborated in Section 2.2. The piecewise structure of $f_{\operatorname{CG}}$ in its parameters is characterized by:
Lemma 3.2 (Lemma 3.2 in Balcan et al. (2021d)).
@@
Proposition 3.3.
Under the same conditions as Theorem 2.5 and 2.6, with the score function $f_{\operatorname{CG}}$,
@@
Lemma 3.4 (Lemma 3.4 in Balcan et al. (2021d)).
@@
Proposition 3.5.
Under the same conditions as Theorem 2.5 and 2.6, with the score function $f_{\operatorname{CG}}^{k}$,
@@ 3.3 Learnability of cutting plane(s) from a finite set
The selection of an optimal cut from an infinite pool of candidate cuts, as discussed in Proposition 3.3 and Proposition 3.5, is often difficult and inefficient in practice. Consequently, a popular way is to select cuts based on information from the simplex tableau (such as GMI cuts), as well as some combinatorial cuts, which inherently limit the number of candidate cuts considered to be finite.
Suppose we have a finite set of cuts $\mathcal{C}$, and we define $f_{\operatorname{ROW}}(I,c)$ as the branch-and-bound tree size after adding a cut $c\in\mathcal{C}$ at the root for a given ILP instance $I\in\mathcal{I}$. Then Corollary 2.7 implies
Proposition 3.6.
Under the same conditions as Theorem 2.5 and 2.6, with the score function $f_{\operatorname{ROW}}$,
$\operatorname{Pdim}\left(\mathcal{F}_{N^{\operatorname{sgn}},\operatorname{Sigmoid}}^{f_{\operatorname{ROW}}}\right)=\mathcal{O}\left(W\log(U|\mathcal{C}|)\right)\text{ and }\operatorname{Pdim}\left(\mathcal{F}_{N^{\operatorname{ReLU}},\operatorname{CReLU}}^{f_{\operatorname{ROW}}}\right)=\mathcal{O}\left(LW\log(U+|\mathcal{C}|)\right).$
3.4 Learnability of cut selection policy
One of the leading open-source solvers, SCIP Gamrath et al. (2020), uses cut selection methodologies that rely on combining several auxiliary score functions. For a finite set of cutting planes $\mathcal{C}$, instead of directly using a neural network for selection, this model selects $c^{*}\in\operatorname{arg\,max}_{c\in\mathcal{C}}\sum_{i=1}^{\ell}\mu_{i}\operatorname{Score}_{i}(c,I)$ for each instance $I\in\mathcal{I}$. Here, $\operatorname{Score}_{i}$ represents different heuristic scoring functions that assess various aspects of a cut, such as "Efficacy" Balas et al. (1996) and "Parallelism" Achterberg (2007), for a specific instance $I$, and the coefficients $\bm{\mu}\in[0,1]^{\ell}$ are tunable weights for these scoring models. Since $\mathcal{C}$ is considered to be finite, the above optimization problem is solved through enumeration. The authors in Turner et al. (2023) have experimentally implemented the idea of using a neural network to map instances to weights. We provide an upper bound on the pseudo-dimension in the following proposition of this learning problem.
Proposition 3.7.
Under the same conditions as Theorem 2.5 and 2.6, let $f_{S}(I,\bm{\mu})$ denote the size of the branch-and-bound tree for $I$ after adding a cutting plane determined by the weighted scoring model parameterized by $\bm{\mu}\in[0,1]^{\ell}$. The pseudo-dimension bounds are given by:
$\operatorname{Pdim}\left(\mathcal{F}_{N^{\operatorname{sgn}},\operatorname{Sigmoid}}^{f_{S}}\right)$ $=\mathcal{O}\left(W\log(U|\mathcal{C}|)\right),$
$\operatorname{Pdim}\left(\mathcal{F}_{N^{\operatorname{ReLU}},\operatorname{CReLU}}^{f_{S}}\right)$ $=\mathcal{O}\left(LW\log(U+\ell)+W\log(|\mathcal{C}|)\right).$
4 Numerical experiments
@@
This problem, concerning the selection from an infinite pool of cuts, presents a significant challenge. The target function $f_{\operatorname{CG}}$, in relation to its parameters, is an intricately complex, high-dimensional, and piecewise constant function with numerous pieces (recall Lemma 3.2), making direct application of classical gradient-based methods seemingly impractical. As such, we use an RL approach, with our goal switched to identify a relatively “good” neural network for this distribution. Then, based on Proposition 3.3 and Theorem 2.3 the average performance of the chosen parameter setting on sampled instances closely approximates the expected performance on the distribution in high probability, given that the sample size $t$ is sufficiently large.
@@ 4.1 Experimental setup
We consider the multiple knapsack problems Kellerer et al. (2004) with 16 items and 2 knapsacks, using the Chvátal distribution as utilized in Balcan et al. (2021b), originally proposed by Chvátal in Chvátal (1980). Our synthetic dataset has a training set of 5,000 instances and a test set of 1,000 instances from the same distribution.
@@
Each instance $I$ sampled from the distribution $\mathcal{D}$ is treated as a state in RL, with the outputs of the neural network considered as actions in the RL framework. The neural network thus functions as the actor in an actor-critic scheme Silver et al. (2014), where the reward is defined as the percentage reduction in the tree size after adding a cut, i.e., $\frac{f_{\operatorname{CG}}(I,\mathbf{0})-f_{\operatorname{CG}}(I,\mathbf{u})}{f_{\operatorname{CG}}(I,\mathbf{0})}$. The Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm Fujimoto et al. (2018) is used here for the training of the neural network.
The experiments were conducted on a Linux machine with a 12-core Intel i7-12700F CPU, 32GB of RAM, and an NVIDIA RTX 3070 GPU with 8GB of VRAM. We used Gurobi 11.0.1 Gurobi Optimization, LLC (2023) to solve the ILPs, with default cuts, heuristics, and presolve settings turned off. The neural networks were implemented using PyTorch 2.3.0. The details of the implementation are available at https://github.com/Hongyu-Cheng/LearnCGusingNN.
@@ 4.2 Empirical results
The experimental results indicate that even a suboptimal neural network parameterization can outperform the cut selection methodologies used in leading solvers like SCIP. Figure 1 presents the average tree size comparison across 1,000 novel test instances, with three distinct strategies:
  1. 1.
    The blue solid line represents the tree size when the cut is selected based on the highest convex combination score of cut efficacy and parallelism, adjusted by a parameter $\mu\in[0,1]$ (incremented in steps of 0.01). The candidate set of cuts includes all CG and GMI cuts generated from the appropriate rows of the optimal simplex tableau.
  2. 2.
    The green dash-dotted line demonstrates a notable reduction in tree size when using cuts generated through our RL approach;
  3. 3.
    The purple dash-dotted line follows the same approach as 2., but uses linear threshold neural networks to generate CG cut parameters. The training process uses the idea of Straight-Through Estimator (STE) Bengio et al. (2013); Yin et al. (2019).
Figure 1: Comparison of branch-and-bound tree sizes using different cut selection strategies.
@@
Table 1: Comparison of the computational speeds between ReLU neural network inference and LP solving. This table presents the total time in seconds for 100 runs on a test set of 1,000 instances.
@@
Computing cuts via trained neural network on GPU 0.010
Computing cuts via trained neural network on CPU 0.055
Solving required LPs using Gurobi 2.031
@@ 5 Discussions and open questions
In our study, we concentrated on adding CG cuts solely at the root of the branch-and-cut tree. However, devising a strategy that generates high quality cutting planes while being efficient across the entire branch-and-cut tree poses a significant and intriguing challenge for future research. Further, our theoretical findings are applicable to any encoder that maps instances into Euclidean spaces. Hence, utilizing a fixed encoder capable of converting ILPs with different number of constraints into a same Euclidean space can in principle enable the training of a unified neural network to generate cutting planes across the branch-and-cut tree. Moreover, an effective encoder could improve the neural network’s performance beyond the one achieved with the basic stacking encoder used in our paper.
The neural network training problem (2) also requires further study. We used an RL approach (see Section 4) to update the neural network parameters and minimize the average tree size for ILP instances sampled from a given distribution. However, this method does not guarantee convergence to optimal parameters, and relies heavily on the exploratory nature of the RL algorithm. For ILP distributions where random exploration of Chvátal multipliers is unlikely to yield smaller tree sizes, the RL algorithm may struggle to identify an effective parameter setting. Developing a more efficient and robust training methodology would greatly improve the practical value of our work.
Acknowledgments and Disclosure of Funding
@@
Appendix A Auxiliary lemmas
Lemma A.1.
For any $x_{1},\dots,x_{n},\lambda_{1},\dots,\lambda_{n}>0$, the following inequalities hold:
$\log x_{1}$ $\leq\frac{x_{1}}{\lambda_{1}}+\log\left(\frac{\lambda_{1}}{e}\right),$ (3)
$x_{1}^{\lambda_{1}}\cdots x_{n}^{\lambda_{n}}$ $\leq\left(\frac{\lambda_{1}x_{1}+\cdots+\lambda_{n}x_{n}}{\lambda_{1}+\cdots+\lambda_{n}}\right)^{\lambda_{1}+\cdots+\lambda_{n}}.$ (4)
Lemma A.2 (Theorem 5.5 in Matousek [1999], Lemma 17 in Bartlett et al. [2019], Lemma 3.3 in Anthony et al. [1999], Theorem 1.3 in Edelsbrunner [1987]).
Let $\mathcal{P}\subseteq\mathbb{R}^{\ell}$ and let $f_{1},\ldots,f_{t}:\mathbb{R}^{\ell}\to\mathbb{R}$ with $t\geq\ell$ be functions that are polynomials of degree $m$ when restricted to $\mathcal{P}$. Then
|$|\{\operatorname{sgn}(f_{1}(\mathbf{p})),\ldots,\operatorname{sgn}(f_{t}(\mathbf{p})):\mathbf{p}\in\mathcal{P}\}|$ $=1,$ $m=0,$
|$|\{\operatorname{sgn}(f_{1}(\mathbf{p})),\ldots,\operatorname{sgn}(f_{t}(\mathbf{p})):\mathbf{p}\in\mathcal{P}\}|$ $\leq 2\left(\frac{2etm}{\ell}\right)^{\ell},$ $m\geq 2.$
Lemma A.3.
$\mathbf{p}\mapsto(h(I_{1},\mathbf{p}),\ldots,h(I_{t},\mathbf{p}))$
Proof.
Within each $\mathcal{P}_{i}$, $h(I_{j},\mathbf{p})-s_{j}$ is a polynomial in $p$ for every $j=1,\ldots,t$, given the hypothesis. Applying Lemma A.2,
Summing over the different $\mathcal{P}_{i}$, $i\in[R(t)]$, we obtain that
Lemma A.4.
Proof.
Follows from Section 2 in Bartlett et al. [1998] and Section 4 in Bartlett et al. [2019]. ∎
Lemma A.5 (Anthony et al. [1999], Sontag et al. [1998]).
Proof.
Fix a $t\in\mathbb{N}$ and $\mathbf{x}^{1},\ldots,\mathbf{x}^{t}\in\mathbb{R}^{w_{0}}$. Consider a neuron in the first hidden layer. The output of this neuron on the input $\mathbf{x}^{j}$ is $\operatorname{sgn}(\langle\mathbf{a},\mathbf{x}^{j}\rangle+b)$, where $\mathbf{a}\in\mathbb{R}^{w_{0}},b\in\mathbb{R}$ are the weights associated with this neuron. As a function of these weights, this is a linear function, i.e., a polynomial function of degree 1. Applying Lemma A.2, there are at most $\left(\frac{et}{w_{0}+1}\right)^{w_{0}+1}$ regions in the space $(\mathbf{a},b)\in\mathbb{R}^{w_{0}}\times\mathbb{R}$ such that within each region, the output of this neuron is constant. Applying the reasoning for the $w_{1}$ neurons in the first layer, we obtain that the output of the first hidden layer (as a $0/1$ vector in $\mathbb{R}^{w_{1}}$) is piecewise constant as a function of the parameters of the first layer, with at most $\left(\frac{et}{w_{0}+1}\right)^{(w_{0}+1)w_{1}}$ pieces. For a fixed output of the first hidden layer (which is a vector in $\mathbb{R}^{w_{1}}$), we can apply the same reasoning and partition space of weights of the second layer into $\left(\frac{et}{w_{1}+1}\right)^{(w_{1}+1)w_{2}}$ regions where the output of the second hidden layer is constant. Applying this argument iteratively across the hidden layers, and using the inequality (4) in Lemma A.1, we deduce that a decomposition exists for $\mathbb{R}^{W}$ with at most
$=$ $\left(\frac{etU}{W^{\prime}}\right)^{W^{\prime}}$
Appendix B Proofs of main results